# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions

### v2

@@ 3.5 Landslide Susceptibility Prediction
Landslide is a geohazard that kill thousands each year. We next evaluated StefaLand for landslide susceptibility prediction using the SLIDO dataset from the State of Oregon, which provides detailed landslide occurrence records. Following Liu et al. (2025), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch. We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
@@
Logistic Regression 0.742 0.707 0.792 0.47 0.819
Random Forest 0.751 0.703 0.834 0.763 0.839
CNN2D 0.778 0.760 0.787 0.773 0.863
StefaLand + CNN2D 0.799 0.791 0.793 0.792 0.875
Note: All baseline results (Logistic Regression, Random Forest, CNN2D) are taken from previously published 30m-resolution experiments in Liu et al. (2025), except for StefaLand + CNN2D, which represents our proposed method.
Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across all metrics except Recall. This is a particularly difficult baseline to improve so even modest gains are rare. ROC AUC increased from 0.863 to 0.875, and precision rose from 0.760 to 0.791, reflecting fewer false positives. While random forest finds a high Recall, this is a particular realization with a large tradeoff on Precision. StefaLand produces overall well-rounded predictions, with both better Precision and Recall than CNN2D.
@@ 4.1 Key Findings and Contributions
As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective (Beery et al., 2018, Gacu et al., 2025). In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models. Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics. These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
The five dynamic prediction cases, along with benchmark models that reproduce state-of-the-art results in the literature, together paint a clear picture. The pretraining of StefaLand on attributes builds deep landscape representations. These features are highly relevant to hydrologic prediction tasks and can avoid overfitting compared to those built in task-specific supervised learning (LSTM or StefaLand-direct), improving model spatial generalization. They are also decidedly more relevant to such tasks than those obtained from existing Earth Foundation models expensively trained on massive amounts of satellite images. It shows that the importance of larger pretraining data may not necessarily exceed that of problem relevance, and image-like data may not be the optimal data representation for such tasks. We stress that our approach is complementary to satellite-based foundation models: where they exploit large-scale visual patterns, StefaLand focuses on problem-relevant attributes, offering an efficient and domain-specific alternative. We have tested a large number of adapter formulations. While we can certainly test more finetuning options, especially internal foundation model layers, the computational and data-storage demands already start to be limiting.
Our attribute-based approach is at least an order of magnitude more efficient than pixel-wise satellite transformers. Our underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining. TerraMind’s larger configuration corresponds to about 7,680 GPU hours (Jakubik et al., 2025), PrithviWxC require roughly 23,040 GPU hours (Schmude et al., 2024) and Aurora required roughly 14,592 GPU hours (Bodnar et al., 2025). In contrast, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on limited budgets. These figures do not include the large gaps in data-storage and transfer demands during pretraining (StefaLand, Terramind and PrithviWxC used $\approx 2$, $\approx 11$ and $\approx 27$ data as mentioned earlier). In fact, it is costly and resource-straining to run the satellite-imagery-focused foundation models, which impedes us from comparing with them in every case. Finally, to the best of our effort, we could not identify other earth foundation models that are designed for land-surface predictions.
@@ Appendix A Detailed Model Architecture
This appendix provides the complete mathematical formulation of the StefaLand model architecture more detailed then 2.
@@ B.1 Physics-Based Differentiable Modeling
For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010, Beck et al., 2020, Bergström, 1976; 1992, Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model. HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables such as evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow. We used an updated modern version, HBV1.1 (Song et al., 2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
@@
Table B1: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
@@ B.2 Linear Regression baselines
Table B2: Additional experiments with linear regression baselines.
@@
Global Streamflow Linear Regression 1.823 1.746 0.252 1.816 1.721 0.248
@@ Appendix C Experimental Details
Table C1: StefaLand Pretraining Configuration
@@
Model MFFormer_dec_LSTM
@@
Table C2: StefaLand Pretraining Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022; 2019)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
Outputs None (self-supervised pretraining)
Table C3: Attribute Groups Used in Group Masking Pretraining
@@
Table C4: CAMELS Streamflow HBV Model Hyperparameters
@@
Table C5: CAMELS Streamflow Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017, Newman et al., 2014)
Outputs Streamflow CAMELS gauge records
Table C6: Soil Moisture Model Configuration
@@
Data loader onlylstm_loader
Data sampler finetuning_noHBV
@@
Table C7: Soil Moisture Variables and Sources
@@
Precipitation Global Precipitation Measurement (GPM), & MSWEP &ERA5 precipitation (Huffman et al., 2019, Beck et al., 2019, Muñoz Sabater, 2019)
Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b; 2011)
Table C8: Global Streamflow Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP (Beck et al., 2022; 2019)
@@
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
@@
Outputs Streamflow Global runoff data (GRDC, 2024)
Table C9: Landslide (SLIDO, Oregon) Variables and Sources
@@
Lithology Global Lithological Map (GLiM) (har)
@@
Table C10: Soil Composition (ISRIC) Variables and Sources
@@
Table C11: PrithviWxC Surface Variables and Sources (MERRA-2)
Variable Type Variable Name Source
Time Series Forcings Precipitation, fluxes, winds, soil moisture/temperature, LAI, runoff, etc. NASA MERRA-2 reanalysis (Gelaro et al., 2017)
Static Attributes Land/ocean/ice fractions, surface geopotential, subgrid orography NASA MERRA-2 constants (Gelaro et al., 2017)
Table C12: Computation Resources for StefaLand and Comparison Models

### v3

@@ 3.5 Landslide Susceptibility Prediction
Landslide is a geohazard that kill thousands each year. We next evaluated StefaLand for landslide susceptibility prediction using the SLIDO dataset from the State of Oregon, which provides detailed landslide occurrence records. Following (Liu et al., 2025a), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch.
@@
Logistic Regression 2D 0.744 0.720 0.795 0.756 0.823
Random Forest 2D 0.765 0.737 0.822 0.777 0.849
CNN2D 0.880 0.896 0.858 0.877 0.954
StefaLand + CNN2D 0.903 0.859 0.963 0.908 0.911
Note: All baseline results (Logistic Regression 2D, Random Forest 2D, CNN2D) are taken from previously published 10m-resolution experiments in (Liu et al., 2025a), except for StefaLand + CNN2D, which represents our proposed method.
We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across most metrics. This is a particularly difficult baseline to improve, so even modest gains are rare. Recall increased from 0.858 to 0.963, and accuracy rose from 0.880 to 0.903, reflecting fewer misclassifications overall. While the random forest achieves high precision in this particular run, it does so at the cost of lower recall. Overall, StefaLand produces well-rounded predictions, with both higher accuracy and recall than CNN2D.
3.6 Ablations
Figure 2: Ablation impact matrix across evaluation settings using RMSE. Each cell shows the percent RMSE increase relative to the full StefaLand ResConn model (lower is better), highlighting the contributions of pretraining and task adaptation. Blank or hatched cells indicate ablations not evaluated for that setting.
To isolate the contributions of pretraining and task adaptation in StefaLand, we conduct a structured ablation study across multiple tasks and spatial evaluation protocols, comparing four variants against the proposed pretrained model with task adapter.
Direct prediction removes pretraining entirely and treats the target as an additional input channel, training the model end to end only on downstream data. Scratch training uses the full StefaLand architecture but trains all parameters from random initialization, isolating the effect of large-scale pretraining. No adapter removes the task-specific adapter and feeds pretrained embeddings directly to the task decoder, testing whether simple feature reuse is sufficient. Linear probing freezes the pretrained encoder and applies only a linear prediction head, representing the weakest form of adaptation.
Across streamflow, soil moisture, and global Caravan experiments, performance degradation relative to the full model is dominated by the removal of pretraining. Direct prediction and scratch training consistently yield the largest errors, especially under spatial generalization. In contrast, removing or simplifying the adapter produces smaller but consistent performance drops, indicating that task adaptation improves results but cannot compensate for the absence of pretrained representations.
These trends are consistent across random and regional holdouts and across tasks with differing data characteristics, highlighting pretrained land-surface representations as the primary driver of StefaLand’s performance. Numerical results for ablations, are provided in Appendix B.4.
@@ 4.1 Key Findings and Contributions
Methods that reliably improve spatial generalization to data-scarce regions remain rare in the literature (Beery et al., 2018; Gacu et al., 2025). In contrast, StefaLand combined with lightweight task-specific heads achieves state-of-the-art or competitive performance across four broad problem classes: streamflow prediction (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models. Across tasks, the strongest gains arise from architectures that combine pretrained StefaLand representations with explicit temporal modeling via residual connections, indicating that attribute-centric pretraining captures problem-relevant structure while downstream sequence models resolve task-specific dynamics. Together, these results support the premise that foundation models can improve out-of-domain transfer and help democratize prediction quality in data-scarce regions.
Across five dynamic prediction settings, evaluated against benchmark models that reproduce state-of-the-art results reported in prior work, a consistent pattern emerges. Pretraining StefaLand on landscape attributes yields deep representations that are highly relevant to hydrologic and land-surface prediction tasks, enabling stronger spatial generalization than purely supervised approaches such as LSTM-based models or training from scratch. Comparisons with alternative pretrained representations further suggest that problem relevance of the pretraining signal is at least as important as scale alone, particularly for tasks governed by physical and environmental processes rather than visual appearance. We emphasize that this attribute-based approach is complementary to satellite-centric foundation models: while image-based models excel at extracting large-scale visual patterns, StefaLand focuses on structured environmental attributes that are directly aligned with land-surface processes, offering a more targeted and efficient alternative.
A key practical advantage of StefaLand is computational efficiency. The attribute-based pretraining strategy avoids pixel-wise processing and large image volumes, resulting in a compact transformer with roughly 12 million parameters and substantially lower data-management requirements. As a result, StefaLand pretraining can be performed on modest computational budgets, making large-scale spatial generalization experiments feasible without access to specialized infrastructure. While additional adapter and finetuning variants could be explored, the current results already demonstrate strong performance across diverse tasks, and further expansion must balance marginal gains against growing computational and storage costs.
@@ Appendix A Detailed Model Architecture
This appendix provides the complete mathematical formulation of the StefaLand model architecture.
@@ B.1 Physics-Based Differentiable Modeling
For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010; Beck et al., 2020; Bergström, 1976, 1992; Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model. HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables such as evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow. We used an updated modern version, HBV1.1 (Song et al., 2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
@@
Table 9: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
@@ B.2 Linear Regression baselines
Table 10: Additional experiments with linear regression baselines.
@@
Caravan Streamflow Linear Regression 2.612 2.431 0.142
@@
B.3 External Foundation Model Comparisons
For completeness, we explored several existing foundation models developed for Earth observation and atmospheric applications, including TerraMind, PrithviWxC, and Galileo (Jakubik et al., 2025; Hsu et al., 2024; Tseng et al., 2025). All models were evaluated using consistent downstream protocols, with pretrained encoders either frozen or minimally adapted and paired with task-specific heads comparable to those used for StefaLand. These experiments were intended to probe the extent to which representations learned from large-scale EO or atmospheric data transfer to land–surface and hydrologic prediction tasks.
These experiments are not intended as exhaustive benchmarks or as performance upper bounds for the evaluated models, but rather as feasibility probes to understand whether their pretrained representations can be directly repurposed for land–surface and geohazard tasks under lightweight adaptation
Because these models differ substantially in their native input formats and pretraining objectives, task-specific adaptations were required. TerraMind and PrithviWxC were coupled with the same residual adaptation architecture used for StefaLand, with only the added adaptation units trained. Due to the intensive data and storage requirements of PrithviWxC, inputs were restricted to surface-level variables most directly related to land–surface interactions, together with static attributes, while multi-level atmospheric variables were excluded. For context, StefaLand, TerraMind, and PrithviWxC were pretrained on approximately 2, 11, and 27 terabytes of data, respectively.
Galileo was evaluated in the landslide susceptibility setting, where inputs consist of multichannel static environmental attributes rather than multispectral time series. To accommodate this difference, we applied an input adaptation strategy that pooled spatial features along one dimension, projected the 17 environmental channels to 12 channels via a linear layer, and treated the remaining spatial dimension as pseudo-temporal input. We initialized Galileo using its pretrained Transformer encoder (768-dimensional embeddings, four attention layers with 12 heads each, and a feed-forward dimension of 3072), followed by a two-layer classification head (768→256→1) with ReLU activation and dropout (rate 0.3). The resulting model was fine-tuned end-to-end using AdamW (learning rate $10^{-4}$, weight decay 0.01), with a batch size of 128 for 1000 epochs on the same train–test split as other landslide experiments, using gradient clipping with a maximum norm of 1.0.
Collectively, these experiments provide a broad exploratory comparison of how foundation models pretrained on EO imagery or atmospheric data behave when adapted to land–surface and geohazard prediction tasks.
Table 11: Performance of external foundation model baselines across tasks. All models use frozen pretrained encoders with the same residual adaptation head.
Task Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
CAMELS Streamflow (PUB) TerraMind-resConn 1.332 $\pm$ 0.0410 1.301 $\pm$ 0.0375 0.777 $\pm$ 0.0071
CAMELS Streamflow (PUR) TerraMind-resConn 1.420 $\pm$ 0.2021 1.398 $\pm$ 0.1932 0.763 $\pm$ 0.0172
Soil Moisture (Random) TerraMind-resConn 0.083 $\pm$ 0.0021 0.062 $\pm$ 0.0007 0.694 $\pm$ 0.0289
PrithviWxC-resConn 0.081 $\pm$ 0.0019 0.060 $\pm$ 0.0004 0.703 $\pm$ 0.0390
Soil Moisture (Europe) TerraMind-resConn 0.101 0.080 0.519
PrithviWxC-resConn 0.103 0.079 0.523
Table 12: Landslide susceptibility prediction using the Galileo foundation model compared with published and proposed baselines on the Oregon SLIDO dataset.
Model (10m) Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ ROC AUC $\uparrow$
Logistic Regression 2D 0.744 0.720 0.795 0.756 0.823
Random Forest 2D 0.765 0.737 0.822 0.777 0.849
CNN2D 0.880 0.896 0.858 0.877 0.854
Galileo + CNN2D 0.750 0.764 0.720 0.742 0.834
StefaLand + CNN2D 0.903 0.859 0.963 0.908 0.911
B.4 Ablations
Table 13: StefaLand ablations on CAMELS streamflow under random (PUB) and regional (PUR) spatial holdout.
Variant PUB: random holdout (ungauged basins) PUR: regional holdout (ungauged regions)
RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
StefaLand direct 1.882 $\pm$ 0.0901 1.849 $\pm$ 0.0691 0.538 $\pm$ 0.0105 0.395 $\pm$ 0.1431 1.982 $\pm$ 0.4012 1.949 $\pm$ 0.3871 0.230 $\pm$ 0.0784 0.201 $\pm$ 1.2201
StefaLand scratch 1.355 $\pm$ 0.0394 1.332 $\pm$ 0.0368 0.801 $\pm$ 0.0031 0.661 $\pm$ 0.0372 1.516 $\pm$ 0.3723 1.378 $\pm$ 0.3496 0.771 $\pm$ 0.0211 0.560 $\pm$ 0.3120
StefaLand noResConn 1.171 $\pm$ 0.0325 1.154 $\pm$ 0.0319 0.823 $\pm$ 0.0022 0.706 $\pm$ 0.3260 1.376 $\pm$ 0.1987 1.356 $\pm$ 0.1712 0.798 $\pm$ 0.0195 0.610 $\pm$ 0.1345
StefaLand linear 1.452 $\pm$ 0.0542 1.366 $\pm$ 0.0533 0.751 $\pm$ 0.0127 0.661 $\pm$ 0.0430 1.484 $\pm$ 0.2588 1.453 $\pm$ 0.2402 0.672 $\pm$ 0.0391 0.542 $\pm$ 0.3100
StefaLand resConn 1.111 $\pm$ 0.0378 1.068 $\pm$ 0.0374 0.869 $\pm$ 0.0067 0.717 $\pm$ 0.1600 1.344 $\pm$ 0.2097 1.334 $\pm$ 0.1873 0.801 $\pm$ 0.0220 0.635 $\pm$ 0.2460
Table 14: StefaLand ablation study on soil moisture prediction under random and regional holdout.
Variant Random location holdout Regional holdout (Europe)
RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
StefaLand direct 0.140 $\pm$ 0.0431 0.103 $\pm$ 0.0041 0.637 $\pm$ 0.0352 0.135 0.112 0.503
StefaLand scratch 0.074 $\pm$ 0.0011 0.058 $\pm$ 0.0003 0.749 $\pm$ 0.0172 0.108 0.064 0.528
StefaLand noResConn 0.075 $\pm$ 0.0009 0.057 $\pm$ 0.0001 0.741 $\pm$ 0.0201 0.095 0.058 0.545
StefaLand linear 0.084 $\pm$ 0.0010 0.061 $\pm$ 0.0002 0.720 $\pm$ 0.0192 0.100 0.063 0.393
StefaLand resConn (proposed) 0.068 $\pm$ 0.0013 0.054 $\pm$ 0.0004 0.783 $\pm$ 0.0054 0.090 0.059 0.638
Figure 3: Adapter ablation on streamflow random spatial splits (PUB). Boxplots summarize per-basin performance distributions for five adapter designs (Feedforward, Bottleneck, MoE, Gated, Residual). We report RMSE (lower is better), correlation (higher is better), and NSE (higher is better) across held-out basins, showing that the Residual adapter yields the most consistent gains, particularly in Corr and NSE.
@@ Appendix C Experimental Details
C.1 Model Configurations and Hyperparameters
Table 15: StefaLand pretraining configuration.
@@ C.1 Model Configurations and Hyperparameters
Model Stefaland_dec_LSTM
@@
Table 16: Attribute groups used in group masking pretraining.
@@
Table 17: CAMELS streamflow HBV model hyperparameters.
@@
Table 18: Soil moisture model configuration.
@@
Data loader onlylstmloader
Data sampler finetuningnoHBV
@@
C.2 Variables and Data Sources
Table 19: StefaLand pretraining variables and sources.
@@ C.2 Variables and Data Sources
Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration from Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022, 2019)
@@
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011; Ramcharan et al., 2018)
@@
Outputs None (self-supervised pretraining)
Table 20: CAMELS streamflow variables and sources.
Variable Type Variable Name Source
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017; Newman et al., 2014)
Outputs Streamflow CAMELS gauge records
Table 21: Soil moisture variables and sources.
Variable Type Variable Name Source
Precipitation Global Precipitation Measurement (GPM), MSWEP, and ERA5 precipitation (Huffman et al., 2019; Beck et al., 2019; Muñoz Sabater, 2019)
Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b, 2011)
Table 22: Streamflow input variables and attributes used from the Caravan dataset.
Variable Type Variable Name Source
Time Series Forcings Precipitation (P), Air temperature (Ta), Potential evapotranspiration (PET) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
Surface pressure, 10 m wind components ($u$, $v$) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
Net solar radiation, net thermal radiation ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
Snow water equivalent, soil moisture (4 layers) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
Static Basin Attributes Latitude, longitude, catchment area Caravan metadata (Kratzert et al., 2023)
Aridity indices (ERA5-Land, FAO Penman–Monteith) Caravan derived attributes (Kratzert et al., 2023)
Mean precipitation, precipitation seasonality Caravan derived attributes (Kratzert et al., 2023)
Mean air temperature, PET seasonality Caravan derived attributes (Kratzert et al., 2023)
Elevation, slope Global terrain products via Caravan (Kratzert et al., 2023)
Soil texture fractions (sand, silt, clay) Harmonized World Soil Database (HWSD) via Caravan (FAO et al., 2012; Kratzert et al., 2023)
Soil erosion index Global soil datasets via Caravan (Kratzert et al., 2023)
Forest cover fraction Global land cover products via Caravan (Kratzert et al., 2023)
Outputs Streamflow Gauge observations compiled in Caravan (Kratzert et al., 2023)
Table 23: Landslide (SLIDO, Oregon) variables and sources.
@@
Lithology Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
@@
Table 24: Soil composition (ISRIC) variables and sources.
@@
C.3 Computational Resources
Table 25: Computation Resources for StefaLand and Comparison Models